Use the following information for the next three (3) questions: In the table below, X represents the number of long distance calls made by a store manager to suppliers during one hour. X 0 1 2 34 19, P(X 2) = (a) 0.2 (b) 0.3 (c) 0.4 (d) 0.7 20. Calculate the expected number of calls (μ) during one hour. (a) μ-10 (b) μ-2 (c) μ = 0.4 (d) μ=0.2
media%2F7d2%2F7d26b9ac-47b9-4e9f-9c14-b6
Answer #1
X P(x) XP(X) (X-\mu _x)^2 (X-\mu _x)^2P(x)
0 0.2 0 4 0.8
1 0.1 0.1 1 0.1
2 0.4 0.8 0 0
3 0.1 0.3 1 0.1
4 0.2 0.8 4 0.8
Mean 2 Variance 1.8

19) Sum of probability= 1

P(0)+P(1)+P(2)+P(3)+P(4)=1

0.2+0.1+P(2)+0.1+0.2=1 =>P(2)= 0.4

20) \mu _x= \sum X_iP(x)= 2

21) \sigma ^2= \sum (X-\mu _x)^2P(x) =1.8